# Testing Quantumness via Photon Statistics for Time-Bin based Quantum Random Number Generators

### v1

@@ Abstract
Randomness is one of the essential components in many fields including cryptography and simulations. Several Quantum Random Number Generator (QRNG) models have been proposed to produce quantum random numbers, which, due to the quantum theory, are more secure than their classical counterparts. However, QRNGs can not produce true random numbers without deterministic classical post-processing. If the underlying distribution of the QRNG is close to a uniform distribution, a small amount of post-processing is sufficient to produce good random numbers retaining quantumness. In this work, we address the randomness and quantumness in the random numbers generated by the QRNGs. We consider two models of QRNGs, which ideally produce random numbers following different distributions (exponential and uniform), and show that, in practice, they are following similar distributions. These empirical photon distributions can be used to test the quantumness of a QRNG. In this letter, we suggest the $\chi^{2}$ goodness-of-fit to test quantumness, as it is known to be an effective method to test if sample data follows a known distribution. We derive a relation when the underlying sampling distributions of the QRNGs will be $\epsilon$-random. Depending on this relation, a suitable post-processing algorithm can be chosen.
@@
Introduction. Random numbers play a crucial role in numerous applications, including computational methods like Monte Carlo simulations and programming Carlson et al. (2015); Metropolis and Ulam (1949), testing principles of physics and chemistry Bell (1964), as well as in the expansive realm of cryptography for creating cryptographic codes or concealing messages Bennett et al. (1992); Shannon (1949), extending to commercial uses such as lottery games and slot machines. In cryptography, randomness is a key input to ensure security in communication. Classical pseudo-random numbers cannot guarantee communication security due to their deterministic generation process. On the other hand, the randomness of the random numbers generated by Quantum Random Number Generators (QRNG) relies on the theory of quantum physics Ma et al. (2016). Different technologies, including shot noise, tunneling, fluctuation, photon arrival time, etc. have been proposed for QRNGs. Many manufacturers are using these technologies to produce their QRNGs. Herrero-Collantes and Garcia-Escartin (2017); ID Quantique (2001); Qutools (2010); Quintessence Labs (2022); Quside (2023); Quantum CTek (2023); Mars Innovation and Department of Electrical and Computer Engineering & Department of Physics, University of Toronto (2023); Quntum Computing Inc. (2023).
@@
Existing time-bin-based QRNG models produce random numbers from exponential Michael A. Wayne and Kwiat (2009); Wayne and Kwiat (2010) and uniform Nie et al. (2014) distributions if the underlying optical devices perform ideally. But, in practice, they cannot be ideal. Therefore, all of the QRNGs require some classical post-processing using deterministic functions. Although such deterministic classical manipulations increase randomness in the generated number, they also reduce quantumness. The amount of classical post-processing should be minimal to retain quantumness in the generated number. For this, the raw number (before classical post-processing) generated by QRNG should be very close to true random.
@@
In this letter, we aim to address photon statistics for practical scenarios, so that we can estimate randomness of the generated raw number. Heavy post-processing should be applied to produce truly random numbers if these raw numbers are less random. Therefore, photon statistics for QRNGs, which we will discuss in this letter, can be considered as one measure of quantumness for photon arrival time-based QRNGs. Depending on this measure, a post-processing algorithm can be chosen to retain the maximum amount of quantumness. Alléaume et al. Alléaume et al. (2004) characterized photon statistics of single-photon sources using the Mandel parameter, and observed that negative values of the Mandel parameter signify non-classical photon statistics for photon sources. Here, we will address the statistics of photons at the detectors. We consider power instability as the error of the source and the attenuators and detection inefficiency as the error of a single photon detector (SPD).
QRNG model. Although producing single photon states is experimentally challenging Imamoḡlu and Yamamoto (1994); Kim et al. (1999), different experimental approaches Brunel et al. (1999); Lounis and Moerner (2000); Santori et al. (2001, 2002); Vučković et al. (2003); Kuhn et al. (2002); Adam and Mechler (2024); Wang and Wang (2024) have been performed to produce single photons. One of the approaches is using a coherent photon source followed by attenuators. This approach produces photon pulses with negligible probability of having multiple photons within a time interval Argillander et al. (2023).
SourceAttenuatorSPDFPGARandom Number
Figure 1: A basic block diagram of the photon arrival time-based QRNG. Coherent pulses, coming from the source, are sent through attenuators to reduce their power so that the probability of having multiple photons within a predefined time interval (typically less than the dead time of SPD) is negligible Argillander et al. (2023). A SPD detects these photons and the detection time (photon arrival time) has been registered. The FPGA provides random numbers based on the registered time stamps.
@@
where $\alpha$ is a complex number and $\hat{a}^{\dagger}$ is creation operator. The mean photon number corresponding to this state is given by $|\alpha|^{2}$. This state can be approximated as a single-photon state with high attenuation. The probability of $n$ photons within $T$ time follows Poisson distribution Titulaer and Glauber (1966) and is given by

### v2

# Quantumness and Randomness of Quantum Random Number Generators via Photon Statistics

@@ Abstract
Randomness is one of the essential components in many fields including cryptography and simulations. Several quantum random number generator (QRNG) models have been proposed to produce random numbers, which, due to the quantum theory, are more secure than their classical counterparts. Many QRNG devices are commercially available as off-the-shelf black-box devices. Computationally, it is not possible to distinguish between a PRNG and a QRNG just by observing their outputs. The common practice for testing quantumness is a direct comparison between the mean and the variance of the experimental photon count. However, statistically, this is not a feasible solution either, because of finite sample size. A QRNG that uses single photons as the quantum source produces true quantum random numbers. Since single photons follow sub-Poissonian statistics, by determining the underlying distribution one can conclude whether a QRNG is truly quantum or not. In this work, we point out the limitations of existing methods of such a decision-making processes, and propose a more efficient two-fold statistical method, which can ensure whether an optical source is quantum or not up to a desired confidence level. Also, QRNGs can not produce true random numbers without deterministic classical post-processing. In this work, we also show that the two models of QRNGs, one producing random numbers from exponential distribution and the other from uniform distribution, become essentially similar under device noise. Detector outputs of both the above models can be tested to quantify the randomness coming from the quantum source of a QRNG, which in turn, dictates how much post-processing is required to produce good random numbers retaining quantumness. In this context, we also derive a relation when the underlying sampling distributions of the QRNGs will be $\epsilon$-random. Depending on this relation, a suitable post-processing algorithm can be chosen.
@@ I Introduction
Random numbers play a crucial role in numerous applications, including computational methods like Monte Carlo simulations and programming Carlson et al. (2015); Metropolis and Ulam (1949), testing principles of physics and chemistry Bell (1964), as well as in the expansive realm of cryptography for creating cryptographic codes or concealing messages Bennett et al. (1992); Shannon (1949), extending to commercial uses such as slot machines and lottery games Konstantinou et al. (2004); Yamanaka and Nakano (2017). In cryptography, randomness is a key input to ensure security in communication. Classical pseudo-random number generators (PRNG) cannot guarantee communication security due to their deterministic generation process. On the other hand, the randomness of the random numbers generated by quantum random number generators (QRNG) relies on the theory of quantum physics Ma et al. (2016). Different technologies, including shot noise, tunneling, fluctuation, photon arrival time, etc. have been proposed for QRNGs. Many manufacturers are using these technologies to produce their QRNGs. Herrero-Collantes and Garcia-Escartin (2017); ID Quantique (2001); Qutools (2010); QNu Labs (2018); Quintessence Labs (2022); Quside (2023); Quantum CTek (2023); Mars Innovation and Department of Electrical and Computer Engineering & Department of Physics, University of Toronto (2023); Quntum Computing Inc. (2023). These QRNGs are commercially available as off-the-shelf black-box devices. Statistically, it is not feasible to distinguish between a PRNG and a QRNG just by observing their outputs. So a pertinent question arises: how do we test whether an alleged QRNG is truly quantum or not?
@@
Let $S_{k}=\{0,1\}^{k}$ be the set of all $k$-bit binary strings. Then $S_{k}$ contains $2^{k}$ elements. A $n$-bit string is called true (or uniformly) random if and only if for any $k$-bit subset ($k\leq n$) $x$ of this string, the probability that $x=y$ for some $y\in S_{k}$ is $\frac{1}{2^{k}}$. Existing time-bin-based QRNG models produce random numbers from exponential Michael A. Wayne and Kwiat (2009); Wayne and Kwiat (2010) or uniform Nie et al. (2014) distributions if the underlying optical devices perform ideally. But, in practice, they cannot be ideal. Therefore, all of the QRNGs require some classical post-processing using deterministic functions. These post-processings use a classical source of randomness to increase randomness in the raw data generated by the QRNGs. Therefore, the amount of classical post-processing should be minimal to ensure the quantum source of the randomness in the final random numbers. For this, the raw number (before classical post-processing) generated by QRNG should be as close to uniformly random as possible.
In this article, we first address the validation of quantum sources using photon counting statistics. Considering that the sample data for the number of photons within a fixed time is available, we suggest two-fold statistical methods to verify whether the source of the sample is quantum or not. Here we assume that the sample only contains the counts detected by the photon detector without any kind of processing. This sample may also contain the dark count of the detector.
Alléaume et al. Alléaume et al. (2004) discussed photon statistics of single-photon sources using the Mandel parameter given by (Mandel, 1979, Equation (11a)). They calculated values of the Mandel parameter from observed photon count timestamps. They concluded that negative values of the Mandel parameter, which indicates sub-Poissonian statistics, signify non-classical photon statistics for photon sources. Some recent experimental works Jones et al. (2023); Peřina et al. (2024a, b); Bush and Børkje (2024) also made the same conclusion. Although analytically it is true, a negative value of the Mandel parameter for experimental data may not always indicate sub-Poissonian statistics. For example, a finite number of photon count timestamps from coherent light that actually follows Poissonian distribution may result in a negative value of the Mandel parameter. This happens because the mean and the variance of a sample with a finite size may not always be the same as the mean and the variance of the underlying distribution. Therefore, calculating the exact value of any parameter from a sample of finite size is not enough to decide the underlying nature of the distribution.
Another approach of testing quantumness of light is to experimentally observe photon antibunching Paul (1982). However, although a stream of antibunched photons indicate the non-classicality nature of light, the converse is not necessarily true. In other words, non-classical light can sometimes be observed as bunched photons Zou and Mandel (1990); Fox (2006). We explain this in more detail in Appendix A.
Our main contribution in this article is to devise a two-fold statistical method to overcome the limitations of the above two existing approaches. The method consists of interval estimation and hypothesis testing, to ensure sub- or super-Poissonian distribution up to a desired confidence level. This can validate the quantumness of light more accurately and efficiently.
@@
As another contribution, we derive the photon statistics for the photon arrival-time-based QRNGs in practical scenarios. We use these statistics to test the randomness of the generated random number. Heavy post-processing should be applied to produce truly random numbers if these raw numbers are less random. Therefore, photon statistics for QRNGs, which we discuss in this article, can be considered as one measure of randomness from quantum sources for photon arrival-time-based QRNGs. Depending on this measure, a suitable post-processing algorithm can be chosen to retain the maximum amount of randomness from the quantum source. For the practical scenario, we consider power instability as the error of the source and the attenuators and detection inefficiency as the error of a single photon detector (SPD).
II Classical vs Quantum Light: Photon-counting Statistics
Photon detection and counting is an important part of QRNG which is further used in other practical applications such as quantum cryptography and quantum communication protocols. In general, photon-based QRNGs generate random numbers by detecting single photons. Therefore, the photon-counting statistics have an important role in determining whether the light falls in the quantum regime or not.
The distribution of the photon-counting statistics can be divided into three classes Fox (2006):
  1. (i)
    sub-Poissonian, i.e., the distribution variance is less than the mean,
  2. (ii)
    Poissonian, i.e., the distribution variance is the same as the mean, and
  3. (iii)
    super-Poissonian, i.e., the distribution variance is greater than the mean.
Classical wave theory of light follows Poissonian and super-Poissonian statistics Fox (2006). On the other hand, the particle nature of light, considered as a sequence of single photons follows sub-Poissonian distributions. Therefore, the sub-Poissonian nature of a photon source is a clear indication of its quantumness Davidovich (1996); Paraïso et al. (2021).
Let us consider a probability distribution given by,
$P(n)=\begin{cases}\frac{e^{-\alpha\mu}(\alpha\mu)^{\alpha n}}{(\alpha n)!},\;n=0,1,2,\dots\,&\text{if }\alpha n\text{ is integer,}\\ 0,&\text{otherwise}\end{cases}$ (1)
for some positive real numbers $\mu,\alpha$. Then the mean would be $\mu$ and variance would be $\frac{\mu}{\alpha}$. Therefore, this denotes a sub-Poissonian, Poissonian, or super-Poissonian distribution for $\alpha>1,\alpha=1,\alpha<1$ respectively.
Thermal light follows Bose–Einstein distribution Fox (2006), given by
$P(n)=\frac{1}{\mu+1}\left(\frac{\mu}{\mu+1}\right)^{n}.$ (2)
The mean of this distribution is given by $\mu$, and the variance is $\mu+\mu^{2}>\mu$. Therefore, this is a super-Poissonian distribution and hence thermal light is an example of classical light.
Next consider a perfect single-photon source that emits photons maintaining equal time interval, $\Delta t$ between two consecutive photons. Then the photon count within time $T$ would be the integer $\mu=\lfloor\frac{T}{\Delta t}\rfloor$. In this case, the mean photon count is given by $\mu$, but the variance is $0<\mu$. Therefore, this is an example of non-classical or quantum light.
@@ II Classical vs Quantum Light: Photon-counting Statistics
where $\alpha$ is a complex number and $\hat{a}^{\dagger}$ is the creation operator. The mean photon number corresponding to this state is given by $|\alpha|^{2}$. This state can be approximated as a single-photon state with high attenuation. The probability of $n$ photons within $T$ time follows Poisson distribution Titulaer and Glauber (1966) given by
@@
where $\mu$ is the mean number of photons per unit time. One can show that for this distribution, the variance of the number of photons per unit time is also $\mu$.
It is interesting to note that mean $=$ variance does not always imply that the distribution is Poisson. For example, consider the probability distribution
$P(k)=(1-p)^{k-1}p,k=1,2,3,\dots,$ (5)
for $0\leq p\leq 1$. If $p=\frac{1}{2}$, this distribution also has mean = variance $=2$. But, it is not Poisson. This motivates us to propose a two-fold method to test the quantumness of the photon source. First, a statistical test is performed to verify whether the mean is equal to or less than or greater than the variance. Note that any statistical validation of equality is not exact, rather approximate in practice. Thus, if the above test validates the mean and the variance to be equal, we propose to perform a second test to validate that the distribution is indeed Poisson.
III Testing quantumness by two-fold Statistical Method
In statistics, an interval $(a,b)$ is called the confidence interval for a distribution parameter $\Theta$, if $P(A<\Theta<B)=1-\epsilon$ for some predefined $\epsilon\in(0,1)$, where $A,B$ are the random variables corresponding to the statistics $a,b$ (that are functions of sample values) respectively. The quantity $1-\epsilon$ is called confidence level of the interval. A statistic $a$ is called an unbiased estimate of an unknown parameter $\Theta$ if the expectation $\langle A\rangle=\Theta$. A statistic $a$ is called a consistent estimate of a parameter $\Theta$ if for every real $\epsilon>0$, $\lim_{n\to\infty}P(|A-\Theta|\geq\epsilon)=0$, where $n$ is the sample size.
Since Poissonian and super-Poissonian statistics are explainable using classical theory, only sub-Poissonian distribution can guarantee the quantumness of a quantum device. We suggest a procedure to decide whether the underlying distribution is sup-Poissonian or not by estimating the confidence interval for the mean of the underlying distribution as follows.
Suppose, the number of photon counts within a fixed time is given as a sample of size $n$. Then by the central limit theorem, $\frac{\langle X\rangle-m}{\sigma/\sqrt{n}}$ is approximately standard normal for large $n$  Hogg and Craig (2019), where $\langle X\rangle$ is the random variable corresponding to the sample mean $\langle x\rangle$, and $m,\sigma^{2}$ are mean and variance of the actual photon-count distribution. Let the sample variance be given by $S^{2}$. Then $s^{2}=\frac{n}{n-1}S^{2}$ is an unbiased and consistent estimate of $\sigma^{2}$ Hogg and Craig (2019). Therefore, by replacing $\sigma$ by $s$, we can say that $Y=\frac{\langle X\rangle-m}{s/\sqrt{n}}$ is approximately standard normal for large $n$.
Let $1-\epsilon$ be the confidence level. Suppose $y_{\epsilon}$ be such that
[…]
Now,
$P(-y_{\epsilon}<Y<y_{\epsilon})$ $=P(-y_{\epsilon}<\frac{\langle X\rangle-m}{s/\sqrt{n}}<y_{\epsilon})$
$=P(\langle X\rangle-\frac{sy_{\epsilon}}{\sqrt{n}}<m<\langle X\rangle+\frac{sy_{\epsilon}}{\sqrt{n}}).$
Therefore, $P(\langle X\rangle-\frac{sy_{\epsilon}}{\sqrt{n}}<m<\langle X\rangle+\frac{sy_{\epsilon}}{\sqrt{n}})=1-\epsilon$. So, we can say that $(\langle x\rangle-\frac{sy_{\epsilon}}{\sqrt{n}},\langle x\rangle+\frac{sy_{\epsilon}}{\sqrt{n}})$ is an approximate confidence interval for the distribution mean $m$ with confidence level $1-\epsilon$.
Now we can decide the photon count distribution as sub-Poissonian, or super-Poissonian, with confidence level $1-\epsilon$, if the estimated variance, $s\leq\langle x\rangle-\frac{sy_{\epsilon}}{\sqrt{n}}$, or $s\geq\langle x\rangle+\frac{sy_{\epsilon}}{\sqrt{n}}$, i.e., $\frac{\langle x\rangle}{s}\geq 1-\frac{y_{\epsilon}}{\sqrt{n}}$, or $\frac{\langle x\rangle}{s}\leq 1+\frac{y_{\epsilon}}{\sqrt{n}}$ respectively, where $y_{\epsilon}$ is given by Equation (6).
Alternatively, one can perform hypothesis testing to determine whether the light is sub-Poissonian, Poissonia, or super-Poissonian.
Hypothesis testing is a statistical procedure to decide whether a statistical hypothesis, called the null hypothesis $H_{0}$, should be rejected or not. A critical region or rejection region of hypothesis testing is a specific region of test statistic values, for which the null hypothesis is rejected. The test statistic for a hypothesis test is a function of the testing sample Casella and Berger (2002). A hypothesis test is called one-tailed if the critical region appears only at one tail of the reference distribution of the test statistic. On the other hand, it is called a two-tailed test, if the critical regions appear at both tails of the above distribution.
During hypothesis testing, two types of error may occur. Type I error occurs when $H_{0}$ is rejected given that it is true. Type II error occurs when $H_{0}$ is accepted given that it is false. Let $\alpha$ and $\beta$ be the probabilities of Type I and Type II errors respectively. The value $\alpha$ denotes the area of the critical region and is also called the significance level. The value $1-\beta$, that is, the probability that $H_{0}$ is rejected given that it is false, is also called the power of a test.
A hypothesis test starts with two hypotheses: the null hypothesis ($H_{0}$) and the alternative hypothesis ($H_{a}$). Depending on the significance level $\alpha$, the critical region is decided. The acceptance or rejection decision of $H_{0}$ is based on what is called the $p$-value of the test statistic. It is defined as $P(|X|>|x|)$, where $X$ is a random variable, whose distribution is the reference distribution of the test statistic of a hypothesis test, and $x$ is the observed value of this statistic Ross (2009); Hogg and Craig (2019). When the observed test statistic falls within the critical region, $p$-value of the test statistic becomes less than $\alpha$. In this case, $H_{0}$ is rejected. If the $p$-value of the test statistic is greater than $\alpha$, the observed test statistic falls outside the critical region. In that case, $H_{0}$ will be accepted.
To determine the quantumness one has to perform two consecutive hypothesis testing. For both the tests $H_{0}:m=s^{2}$. However, the alternative hypotheses $H_{a}$ would be $m<s^{2}$, and $m>s^{2}$ for the first and second test respectively. For both tests, the test statistic would be $z=\frac{\langle x\rangle-s^{2}}{s/\sqrt{n}}$. For the first test, if $z\geq y_{\epsilon}$, where $y_{\epsilon}$ is given by Equation (6), $H_{0}$ is rejected. That implies the underlying photon distribution is sub-Poissonian. But if $z<y_{\epsilon}$, we cannot reject $H_{0}$. Also, we cannot accept $H_{0}$ due to lack of evidence. In that case, we have to perform second hypothesis testing with the same statistic. If $z\geq y_{\epsilon}$, $H_{0}$ is rejected again, and we can say that the underlying distribution is super-Poissonian. Note that, this hypothesis testing is equivalent to the above interval estimation technique, and both results in the same decision.
As we saw in Section II, the equality of mean and variance may not imply Poisson distribution. Therefore to ensure the Poisson distribution which is an indicator of a coherent source of light, we suggest a second test which is a goodness-of-fit test, that is, a hypothesis testing to decide whether a sample data comes from a Poisson distribution or not. Chi-square ($\chi^{2}$) testCochran (1952), Kolmogorov–Smirnov test (KS test) Massey (1951), and Anderson–Darling test Anderson and Darling (1952) are some examples of the goodness-of-fit test Dodge (2008a). Out of these tests, the last two tests are designed for continuous distributions only. However, the KS test is modified later to consider discrete distributions as well Conover (1972). The $\chi^{2}$ test has been found more robust than the KS test for the continuous distributions Slakter (1965).
Figure 1: Probability density function for $\chi^{2}$ distribution. The critical region ($\chi^{2}>\chi^{2}_{k,\alpha}$) for the $\chi^{2}$ test is denoted by stars, and the region corresponding to the $p$-value ($\chi^{2}>\chi^{2}_{obs}$) is denoted by dots. Since here $\chi^{2}_{k,\alpha}>\chi^{2}_{obs}$, the $p$-value becomes larger than $\alpha$ and the region corresponding to the $p$-value contains the critical region. Therefore, in this case, the null hypothesis is accepted.
For testing the quantumness in the generated random number, one may use the $\chi^{2}$ goodness-of-fit test Hogg and Craig (2019); Dodge (2008b) on the data generated by QRNGs as follows. Consider null and alternative hypotheses as
H0: The data follow the distribution (15).
[…]
The test statistic is defined as
$\chi^{2}=\sum_{n=1}^{N}\frac{(O_{n}-E_{n})^{2}}{E_{n}},$ (7)
where $O_{n}$ denotes the frequency of $n$ number of photon in the data, $E_{n}$ denotes the frequency given by Poisson distribution (4). This test statistic follows $\chi^{2}$ distribution with $k=N-1$ degrees of freedom if all the parameters are known, and its density function is given by,
$f_{\chi^{2}}(x)$ $=\frac{x^{\frac{k}{2}-1}e^{-\frac{x}{2}}}{2^{\frac{k}{2}}\Gamma(\frac{k}{2})},\quad x\geq 0,$ (8)
[…]
is the well known Gamma function Hogg and Craig (2019). The cumulative distribution function in this case is given by,
[…]
Figure 1 is a plot of the probability density function of a $\chi^{2}$ distribution. The critical region ($\chi^{2}>\chi^{2}_{k,\alpha}$) for the $\chi^{2}$ test is denoted by stars, and the region corresponding to the $p$-value ($\chi^{2}>\chi^{2}_{obs}$) is denoted by dots. The area of the critical region is given by $P(\chi^{2}>\chi^{2}_{k,\alpha})=1-F_{\chi^{2}}(\chi^{2}_{k,\alpha})=\alpha$. $\chi^{2}$-test being a one-tailed test, $p$-value for this test is given by $P(\chi^{2}>\chi^{2}_{obs})=1-F_{\chi^{2}}(\chi^{2}_{obs})$, where $\chi^{2}_{obs}$ is the observed value of the test statistic.
If the typical values of the parameters $\mu,T$ in Equation (4) are known, we have $k=N-1$. However, if they are unknown, in that case, $k=N-2$. Therefore, $H_{0}$ is rejected if $\chi^{2}_{obs}>\chi^{2}_{k,\alpha}$, where $\chi^{2}_{k,\alpha}$ is the value given by $F_{\chi^{2}}(\chi^{2}_{k,\alpha})=1-\alpha$. In other words, $H_{0}$ is rejected if $p$-value is less than $\alpha$.
As mentioned in Section II, in most cases, the equality of the mean and the variance denotes Poisson distribution, for practical purposes, the last $\chi^{2}$ test may not be needed. We mention this test for the completeness of the discussion.
IV Testing Quantum Randomness of QRNG: Ideal vs Practical
IV.1 QRNG based on single-photon source
QRNG, based on single photons, uses $/50$ 50 beam splitter Jennewein et al. (2000), where an $R/T$ beam-splitter denotes an optical device that splits an incoming light beam into $R\%$ reflected and $T\%$ transmitted light beams. A standard technique is to place a $/50$ 50 Beam-splitter making an angle of $45^{\circ}$ with the path of the photons. Thus, a single photon either passes through the beam-splitter, or it is reflected with probability $\frac{1}{2}$. Two photon detectors say $D_{0}$ and $D_{1}$, can be placed over two output paths. If a photon is detected by $D_{i}$, $i$ is registered as a random bit. A block diagram of this process is given in Figure 2.
[…]
Figure 2: A generic block diagram of the single-photon-based QRNG. Single-photon, coming from the source, are sent through a $/50$ 50 beam-splitter. Two SPDs ($D_{0},D_{1}$) are placed in two paths to detect the photon. Photon may take either path with probability $0.5$. The random bit would be $i$ if the detection happens at $D_{i}$.
If the photon source for this QRNG is not a perfect single-photon source, it may happen that at the same time some photons are transmitted through the beam-splitter and some are reflected. Then, both the detectors may detect photons simultaneously, and the random number cannot be generated. To ensure whether the source is a single-photon source or not, one may perform the first step of the two-fold statistical method mentioned in Section III. If this method reveals that the distribution is not sub-Poissonian, then the source cannot be a single-photon source, and that source cannot be used in this QRNG.
Each bit of the random number generated with this single-photon-based QRNG has probability $\frac{1}{2}$ of being 00 or $1$. So, ideally it produces uniform random numbers. However, any defect in the beam-splitter or the detectors may change this probability. For example, consider an $R/T$ beam-splitter and the detectors $D_{i}$ with detection efficiency $d_{i}\in(0,1]$. Suppose, $D_{0}$ is placed on the path of the transmitted light, and $D_{1}$ is placed on the path of the reflected light. Then the probability of 00 in the generated random number, that is, the probability of detecting a photon by detector $D_{0}$ would be $\frac{d_{0}}{d_{0}+d_{1}}\frac{T}{100}$. Similarly, the probability of $1$ would be $\frac{d_{1}}{d_{0}+d_{1}}\frac{R}{100}$. If $d_{0}T\neq d_{1}R$, the generated random number would not be uniformly random. To test this randomness, statistical tests like NIST Bassham et al. (2010), Dieharder Brown et al. (2014), ENT Walker (2008) etc. may be used. See Appendix B for a discussion on these.
IV.2 QRNG based on Coherent Pulse
Producing single photon states is experimentally challenging Imamoḡlu and Yamamoto (1994); Kim et al. (1999). However, different experimental approaches Brunel et al. (1999); Lounis and Moerner (2000); Santori et al. (2001, 2002); Vučković et al. (2003); Kuhn et al. (2002); Adam and Mechler (2024); Wang and Wang (2024) have been performed to produce single photons. One of the approaches is using a coherent photon source followed by attenuators. This approach produces photon pulses with negligible probability of having multiple photons within a time interval Argillander et al. (2023). In this process, the photons follow Poissonian statistics.